OpenStudy (anonymous):

What is the maximum number of turns in the graph of f(x)=2x³-2x²+7x-25(?)

jimthompson5910 (jim_thompson5910):

In general, if the degree is n, then there are n-1 maximum turns on the graph. In the case of f(x) = 2x³-2x²+7x-25, the degree is n = 3 and n-1= 3-1 = 2 So at most, there are 2 turns on the graph.

OpenStudy (anonymous):

 Cubic hold this general patern